OpenStudy (anonymous):

Solve for x. 7(x - 3) = 4(x + 5)

OpenStudy (anonymous):



OpenStudy (anonymous):

x = 1/11 x = 1/3 x = 41/11 x = 41/3

OpenStudy (anonymous):

those are the chocies I have to chose from